OpenStudy (anonymous):

how do i find the zeros of the polynomial function, f(x)=2x^3-9x+3? pleassseee help!

OpenStudy (bahrom7893):

set that equal to 0

OpenStudy (bahrom7893):

2x^3 - 9x + 3 = 0

OpenStudy (bahrom7893):

now try factoring using long division.. lemme see

OpenStudy (anonymous):

okay, so frist what i do is set f(x) to 0? and solve?

OpenStudy (bahrom7893):

yeah

OpenStudy (anonymous):

I think it shoild be 2x^3-9x+2

OpenStudy (anonymous):

Are you sure it's raised to the 3rd power and not the second?

OpenStudy (anonymous):

no dipin, it is +3 not +2 veronometry, yes it is rasied to the 3rd degree

OpenStudy (bahrom7893):

im trying to factor using long division lol..

OpenStudy (anonymous):

what i don't understand is the steps i need to find the 0's?

OpenStudy (bahrom7893):

wait im working on it..

OpenStudy (anonymous):

okay thanks bahrom7893

OpenStudy (bahrom7893):

u can't solve here.. i just tried factoring with wolf..

OpenStudy (anonymous):

what i need help with more than anything is understanding what i need to do in order. i suppose like a step by step list to understand how to find the zeros of a polynomial function

OpenStudy (anonymous):

what is wolf?

OpenStudy (bahrom7893):

wolframalpha.com

OpenStudy (bahrom7893):

To do step by step, first set the polynomial equal to 0, then try factoring or use the quadratics..

OpenStudy (anonymous):

The zeros of any function is the same as finding the x-intercept. And, since ANY x-intercept has a y value of 0, you set he entire function equal to zero in order to solve for x. That way, when you solve for x, you get ordered pair(s) of (x,0) etc (depending on the degree of your function. There are some exceptions to that rule, but for the most part it's the same. Does that help?

OpenStudy (anonymous):

I undertstand now that i frist need to set the equation equal to 0. other than that do i just try to balance out the equation?

OpenStudy (anonymous):

or do all the x's in the equation turn to 0\

OpenStudy (bahrom7893):

no u just need to find xs that make the y or f(x) go to 0

OpenStudy (anonymous):

okay this is where i have gotten \[2x^3-9x+3=0\] i added 3 to both sides \[2x^3-9x=-3\] i divided each side by 2 \[x^3-9x=-1.5\] now what?

OpenStudy (bahrom7893):

see I dunno if there was a 0 instead of -1.5 we could solve it... but even if u add a 1.5 to both sides ull end up with a 1.5 on the right..sorry i can't solve it..

OpenStudy (anonymous):

its okay m just trying to find the 0;s not slove the equations

OpenStudy (anonymous):

Bahrom is right, it's unfactorable. That means you have no zeros. Which means there are no x-intercepts. The graph is probably floating above or below the x-axis.

OpenStudy (anonymous):

Btw, you can't divide both sides by 2. You can only do that if you factor out a 2 from both of the terms.

OpenStudy (anonymous):

okay i believe you because i have no idea what the hell i am doing lol